A number of chemical reactions have been proposed for the formation of organic (chemicals of life) compounds from inorganic precursors. Here is one of them: H2S(g) + FeS(s) + CO2(g) → FeS2(s) + HCO2H(l)

a) identify the ions and charges in FeS and FeS2 and name the compounds

b) How much HCO2H is obtained by reacting 1.10 g of FeS with 0.500 g H2S and 0.500 g of CO2 if the reaction produces a yield of 50%?  
Answer #1

a) Fe2+ , S2-, and S22-

The ion names would be iron(II), sulfide, and disulfide, and you would just combine them to make the names of the iron compounds.

FeS = iron (II) sulfide
FeS2 = iron (II) persulfide

The other compounds are hydrogen sulfide and formic acid.

b) (1.10g FeS)/(87.910g FeS/mol FeS) = 0.0125mol FeS

(0.5g H2S)/(34.081g H2S/mol H2S) = 0.0147mol H2S

There's a slight excess H2S. Hence FeS is the limiting reactant.

Theoretically we would get

(1.10g FeS)/(87.910g FeS/mol FeS)(1mol HCO2H/1mol FeS)(46.025g HCO2H/mol HCO2H)

With 50.0% yield, you will get

(1.10g FeS)/(87.910g FeS/mol)(1mol HCO2H/1mol FeS)(46.025g HCO2H/mol)(0.50) = 0.2879g HCO2H